You want to buy a bonus that will have a redemption value of $ 20,000. The bonus pays 6% per year paid twice a year. The bonus is redeemable in 15 years. If you want to earn 12% compound annual interest per month, how much would you be willing to pay for the bonus today?
Answer #1

The redemption value of bonus = $20000

Interest rate(r) = 6% per year paid twice in a year, compounding period = 2

Redeemable period = 15 years

Therefore, bonus amount at 6% interest rate after 15 years = 20000(1+.06/2)15.2 = $48545

Let you are willing to pay $p for the bonus today which amounts to $48545 after 15 years which is redeemable period. Compound interest rate = 12% per year, compounding period =12

Then $48545 = $P(1+12/1200)15.12 or P= 48545/(1+0.01)180 = $8096 which is amount you will be willing to pay for bonus today.